### v2

# Taylor-Accelerated Neural Network Interpolation Operators on Irregular Grids with Higher Order Approximation

Abstract
In this paper, a new class of Taylor-accelerated neural network interpolation operators is introduced on quasi-uniform irregular grids. These operators improve existing neural network interpolation operators by incorporating Taylor polynomials at the sampling nodes, thereby exploiting higher smoothness of the target function. The proposed operators are shown to be well defined, uniformly bounded, and to satisfy an exact interpolation property at the grid points. In addition, polynomial reproduction up to a prescribed degree is established. Jackson-type approximation estimates are derived in terms of higher-order moduli of smoothness, yielding enhanced convergence rates for sufficiently smooth functions. Numerical experiments are presented to support the theoretical analysis and to demonstrate the significant accuracy improvement achieved through the Taylor-accelerated construction. In particular, higher-order convergence on irregular grids is obtained, and the proposed approach outperforms existing neural network interpolation operators on irregular grids, including Lagrange-based schemes.
Keyword: Sigmoidal function, NNOs, Taylor approximation, Uniform approximation, Jackson-type estimates.
MSC 2020 41A25, 41A30, 41A46, 41A58, 47A58, 92B20.
1 Introduction and Preliminaries
Neural networks are widely recognized for their remarkable universal approximation capabilities and have become fundamental tools in applied mathematics, scientific computing, and numerical analysis. Beyond the classical training-based paradigm, the theory of neural network operators (NNOs) provides a constructive approximation framework in which neural network architectures are interpreted as linear operators acting on function spaces. This operator-based viewpoint allows one to derive explicit convergence theorems together with quantitative error bounds in the spirit of classical approximation theory.
The systematic study of NNOs was initiated by Cardaliaguet and Euvrard [4], who introduced bell-shaped and squashing-type NNOs and established early connections between neural network structures and polynomial-type approximation processes. Subsequently, Anastassiou and his collaborators [1, 2, 3] developed a broad approximation theory for sigmoidal-based neural network operators, including multivariate extensions and convergence rates in various normed settings. Afterward, several authors investigated NNOs using different activation functions and normed spaces for the functions under approximation [8, 11, 5, 6, 18].
On the other hand, existing NNOs often achieve only low-order approximation, since their error is typically governed by the first modulus of continuity even when the target function possesses additional smoothness. This motivates the problem of designing neural network operator capable of exploiting higher regularity and attaining improved convergence orders. To address this limitation, Kadak[14] recently proposed Hermite-type NNOs incorporating derivative information of the approximated function on uniform grids.
In recent years, increasing attention has been devoted to interpolation-type NNOs, which aim to reconstruct functions directly from sampled data. One of the first constructive interpolation operators within the NNOs framework was proposed by Costarelli [9], and later extended to multivariate settings in [10]. Further developments include interpolation operators activated by smooth ramp functions, providing quantitative rates in terms of moduli of continuity and smoothness (see, e.g., [17, 20, 15, 16]).
However, most existing neural network interpolation constructions are developed on uniform grids, whereas in many practical applications sampling nodes arise naturally in an irregular and non-uniform manner. Addressing this issue, Costarelli et.al. [7], recently introduced a new family of neural network interpolation operators on irregular grids, based on Lagrange polynomials and higher-order moduli of smoothness.
Inspired by higher-order sampling methods and local polynomial approximation techniques, the present work introduces a new class of Taylor-accelerated neural network interpolation operators on irregular grids. The main idea is to enrich sampling-based neural network interpolation operators by replacing pointwise function values with local Taylor polynomials of degree $r$, thereby incorporating derivative information into the approximation mechanism. As a result, the proposed operators preserve exact interpolation at the grid nodes while achieving approximation order $r+1$ for sufficiently smooth functions, with error estimates expressed in terms of higher-order moduli of smoothness.
Now, we collect some definitions and tools that will be used throughout the paper. In particular, we recall the functional setting, the notion of irregular grids, the class of compactly supported sigmoidal activation, and higher-order moduli of smoothness.
Throughout the paper, we consider real-valued functions defined on the compact interval $c,d]\subset\mathbb{R}$.
Space of continuous functions. We denote by $C[c,d]$ the Banach space of all continuous functions $g:[c,d]\to\mathbb{R}$, endowed with the uniform norm
$\|g\|_{\infty}:=\max_{y\in[c,d]}|g(y)|.$
Space of smooth functions. For an integer $r\geq 1$, we denote by $C^{r}[c,d]$ the space of all functions whose derivatives up to order $r$ exist and are continuous on $[c,d]$, namely,
$g^{(j)}\ \text{exists and is continuous on }[c,d],\qquad j=0,1,\dots,r.$
Equipped with the norm
$\|g\|_{C^{r}}:=\sum_{j=0}^{r}\|g^{(j)}\|_{\infty},$
the space $C^{r}[c,d]$ is a Banach space continuously embedded into $C[c,d]$.
Since the Taylor-accelerated neural network interpolation operators introduced in this work employ local Taylor polynomials of degree $r$, we shall mainly assume $g\in C^{r}[c,d]$. Moreover, in the derivation of quantitative error estimates we will require the additional smoothness condition $g\in C^{r+1}[c,d]$.
Irregular grids. Let $c=z_{0}<z_{1}<\cdots<z_{n}=d$ be a partition of the interval $[c,d]$. In contrast to uniform grids, the nodes $\{z_{k}\}_{k=0}^{n}$ are allowed to be distributed in a non-uniform manner.
We assume that the mesh sizes satisfy the controlled irregularity condition
$0<d_{1}\leq z_{k+1}-z_{k}\leq d_{2},\qquad k=0,1,\dots,n-1,$
together with the quasi-uniformity condition
$d_{2}<2d_{1}.$
This hypothesis guarantees that the irregular grid does not exhibit excessively large variations in the step sizes and is crucial for the exact interpolation property of the proposed operators (cf. [7]).
Definition 1.1.
A measurable function $\eta:\mathbb{R}\to\mathbb{R}$ is called sigmoidal if
$\lim_{y\to-\infty}\eta(y)=0,\qquad\lim_{y\to+\infty}\eta(y)=1.$
In this work, we restrict ourselves to sigmoidal functions belonging to the class $D(m)$, $m\in\mathbb{R}^{+}$, defined as follows: a sigmoidal function $\eta:\mathbb{R}\to\mathbb{R}$ belongs to $D(m)$ if
  1. 1.
    $\eta$ is increasing;
  2. 2.
    $\eta(y)=0$ for $y\leq-m$ and $\eta(y)=1$ for $y\geq m$;
  3. 3.
    $0<\eta(y)<1$ for m<y<m-m<y<m.
Associated with $\eta\in D(m)$, we define the activation function $\sigma:\mathbb{R}\to\mathbb{R}$ by
$\sigma(y):=\eta(y+m)-\eta(y-m).$
The activation function $\sigma$ satisfies the following fundamental properties:
  • (M1)
    $\sigma(y)\geq 0$ for all $y\in\mathbb{R}$.
  • (M2)
    $\sigma$ is increasing on $(-\infty,0]$ and decreasing on $0,\infty)$.
  • (M3)
    $\sigma$ has compact support, namely
    $\sigma(y)=0,\qquad|y|\geq 2m.$
  • (M4)
    $\sigma(y)+\sigma(y-2m)=1,\qquad y\in[0,2m].$
  • (M5)
    $\sigma(0)=1,\qquad\sigma\!\left(\frac{2m}{d_{1}}(z_{j}-z_{k})\right)=0,\quad j\neq k.$
    Indeed, since the grids $\{z_{k}\}_{k=0}^{n}$ is strictly increasing and satisfies $z_{k+1}-z_{k}\geq d_{1}$ for all $k$, it follows that
    $|z_{j}-z_{k}|\geq d_{1},\qquad j\neq k.$
    Consequently,
    $\left|\frac{2m}{d_{1}}(z_{j}-z_{k})\right|\geq\frac{2m}{d_{1}}\,|z_{j}-z_{k}|\geq 2m.$
    Hence, by the compact support property (M3), we obtain
    $\sigma\!\left(\frac{2m}{d_{1}}(z_{j}-z_{k})\right)=0,\qquad j\neq k.$
For further details on sigmoidal functions of class $D(m)$ and their properties, we refer the reader to [7].
Definition 1.2.
Let $g\in C[c,d]$ and $\delta>0$. The $(r+1)$-th modulus of smoothness of $g$ is defined by
$\omega_{r+1}(g,\delta):=\sup_{|h|\leq\delta}\|\Delta_{h}^{\,r+1}g\|_{\infty},$
where $\Delta_{h}^{\,r+1}$ denotes the forward difference operator of order $r+1$, taken whenever all shifted points remain in $[c,d]$.
Moreover, it is well known that if $g\in C^{r+1}[c,d]$, then
$\omega_{r+1}(g,\delta)\leq C_{r}\,\delta^{r+1}\|g^{(r+1)}\|_{\infty},$
where $C_{r}>0$ depends only on $r$ (see, e.g., [12, 19, 13]).
2 Taylor-Accelerated Neural Network Interpolation Operator on Irregular Grids
In this section, we introduce a Taylor-type neural network interpolation operator constructed on irregular grids. The main idea is to enhance a neural network interpolation operator by incorporating higher-order Taylor information at the grid nodes, while preserving the exact nodal interpolation property.
Let $g\in C^{r}[c,d]$, where $r\geq 1$ is a fixed integer. For each node $z_{k}$ of the irregular grid, we consider the Taylor polynomial of degree $r$ centred at $z_{k}$, defined by
$P_{r}(g;y,z_{k}):=\sum_{j=0}^{r}\frac{g^{(j)}(z_{k})}{j!}(y-z_{k})^{j},\qquad y\in[c,d].$
Definition 2.1.
Let $g\in C^{r}[c,d]$. Then Taylor-accelerated neural network interpolation operator is defined by
$\widetilde{T}_{n,r}(g,y)=\frac{\displaystyle\sum_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)P_{r}(g;y,z_{k})}{\displaystyle\sum_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)},\qquad y\in[c,d],$
where $\sigma$ is the compactly supported activation function introduced in Section 1 and $d_{1}$ denotes the minimum mesh size of the irregular grids.
Now, first we will show the well-definedness and boundedness of the operator $\widetilde{T}_{n,r}(\cdot)$.
Proposition 2.2.
For every $g\in C^{r}[c,d]$ and every $y\in[c,d]$, the operator $\widetilde{T}_{n,r}(g,y)$ is well defined.
Proof.
Consider the denominator
$D(y):=\sum_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right),\qquad y\in[c,d].$
By the compact support property (M3) of $\sigma$, only finitely many terms in the above sum are nonzero for each fixed $y\in[c,d]$.
Fix $y\in[c,d]$. Since $\{z_{k}\}_{k=0}^{n}$ is a partition of $[c,d]$, there exists an index $i\in\{0,\dots,n-1\}$ such that $y\in[z_{i},z_{i+1}]$. Consequently,
$\min\{|y-z_{i}|,|y-z_{i+1}|\}\leq\frac{z_{i+1}-z_{i}}{2}\leq\frac{d_{2}}{2}.$
Let $k_{0}\in\{i,i+1\}$ be such that
$|y-z_{k_{0}}|\leq\frac{d_{2}}{2}.$
Using the mesh regularity condition and the assumption $d_{2}<2d_{1}$, we obtain
$\left|\frac{2m}{d_{1}}(y-z_{k_{0}})\right|\leq\frac{md_{2}}{d_{1}}<2m.$
Hence,
$\frac{2m}{d_{1}}(y-z_{k_{0}})\in(-2m,2m).$
Since $\eta\in D(m)$ satisfies $0<\eta(y)<1$ for m<y<m-m<y<m, it follows that $\sigma(y)>0$ for all $y\in(-2m,2m)$. Therefore,
$\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k_{0}})\right)>0.$
Consequently,
$D(y)\geq\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k_{0}})\right)>0,\qquad y\in[c,d].$
Thus, the denominator in the definition of $\widetilde{T}_{n,r}(g,y)$ is strictly positive for all $y\in[c,d]$, and the operator $\widetilde{T}_{n,r}(g,y)$ is well defined. ∎
Theorem 2.3.
The operator $\widetilde{T}_{n,r}$ is uniformly bounded on $C^{r}[c,d]$. More precisely, there exists a constant $C_{r}>0$, depending only on $r$, the activation function $\sigma$, and the mesh parameters $d_{1},d_{2}$, such that
$\|\widetilde{T}_{n,r}(g)\|_{\infty}\leq C_{r}\,\|g\|_{C^{r}},\qquad g\in C^{r}[c,d].$
Proof.
Let $g\in C^{r}[c,d]$ and fix $y\in[c,d]$. By the definition of $\widetilde{T}_{n,r}$ and the non-negativity of $\sigma$, we have
$|\widetilde{T}_{n,r}(g,y)|\leq\frac{\sum\limits_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)\,|P_{r}(g;y,z_{k})|}{\sum\limits_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)}.$ (2.1)
Recalling that
$P_{r}(g;y,z_{k})=\sum_{j=0}^{r}\frac{g^{(j)}(z_{k})}{j!}(y-z_{k})^{j},$
we obtain
$|P_{r}(g;y,z_{k})|\leq\sum_{j=0}^{r}\frac{\|g^{(j)}\|_{\infty}}{j!}\,|y-z_{k}|^{j}.$ (2.2)
By the compact support property (M3) of $\sigma$, if
$\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)\neq 0,$
then
$|y-z_{k}|<d_{1}.$
So, for all $j=0,1,\dots,r$,
$|y-z_{k}|^{j}\leq d_{1}^{\,j}.$
Combining this with (2.2), we obtain
$|P_{r}(g;y,z_{k})|\leq\left(\sum_{j=0}^{r}\frac{d_{1}^{\,j}}{j!}\right)\|g\|_{C^{r}}=:M_{r}\,\|g\|_{C^{r}}.$
Substituting into (2.1) yields
$|\widetilde{T}_{n,r}(g,y)|\leq\frac{M_{r}}{\sum\limits_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)}\|g\|_{C^{r}}.$
By Proposition 2.2, there exists a constant
$c_{\sigma}:=\sigma\!\left(\frac{md_{2}}{d_{1}}\right)>0$
such that
$\sum_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)\geq c_{\sigma},\qquad y\in[c,d].$
Therefore,
$|\widetilde{T}_{n,r}(g,y)|\leq\frac{M_{r}}{c_{\sigma}}\,\|g\|_{C^{r}},\qquad y\in[c,d].$
Taking the supremum over $y$ completes the proof. ∎
Remark 2.4.
The uniform boundedness of $\widetilde{T}_{n,r}$ ensures the stability of the Taylor-accelerated neural network interpolation operator. In particular, small perturbations of $g$ in the $C^{r}$-norm lead to controlled perturbations in the approximation $\widetilde{T}_{n,r}(g)$.
2.1 Approximation Estimate
In this subsection, we study the approximation properties of the Taylor-accelerated neural network interpolation operator $\widetilde{T}_{n,r}$. In particular, we establish its polynomial reproduction and exact interpolation properties, and derive Jackson-type approximation estimates in terms of the modulus of smoothness. Several corollaries are also presented to illustrate the convergence behavior under additional regularity assumptions.
Proposition 2.5 (Polynomial Reproduction).
Let $r\geq 1$ and let $p$ be a polynomial of degree at most $r$. Then the Taylor-accelerated neural network interpolation operator reproduces $p$ exactly, namely,
$\widetilde{T}_{n,r}(p,y)=p(y),\qquad\forall\,y\in[c,d].$
Proof.
Since $\deg(p)\leq r$, we have $p^{(r+1)}\equiv 0$. Therefore, the Taylor polynomial of degree $r$ of $p$ about any node $z_{k}$ coincides with $p$ itself, i.e., for all $y\in[c,d]$ and all $k=0,1,\dots,n$,
$P_{r}(p;y,z_{k})=p(y).$
Substituting this identity into the definition of $\widetilde{T}_{n,r}$, we obtain
$\widetilde{T}_{n,r}(p,y)=\frac{\displaystyle\sum_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)p(y)}{\displaystyle\sum_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)}.$
Since $p(y)$ is independent of $k$, it can be factored out, yielding
$\widetilde{T}_{n,r}(p,y)=p(y)\,\frac{\displaystyle\sum_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)}{\displaystyle\sum_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)}=p(y).$
This proves the polynomial reproduction property. ∎
Theorem 2.6 (Interpolation Property).
Let $g\in C^{r}[c,d]$. Then, for every grid node $z_{j}$, $j=0,1,\dots,n$, the Taylor-accelerated neural network interpolation operator satisfies
$\widetilde{T}_{n,r}(g,z_{j})=g(z_{j}).$
Proof.
Fix a node $z_{j}$. By definition,
$\widetilde{T}_{n,r}(g,z_{j})=\frac{\displaystyle\sum_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(z_{j}-z_{k})\right)P_{r}(g;z_{j},z_{k})}{\displaystyle\sum_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(z_{j}-z_{k})\right)}.$
First, the Taylor polynomial of degree $r$ centered at $z_{j}$ satisfies
$P_{r}(g;z_{j},z_{j})=\sum_{s=0}^{r}\frac{g^{(s)}(z_{j})}{s!}(z_{j}-z_{j})^{s}=g(z_{j}).$
Next, if $k\neq j$, then since the grids $\{z_{k}\}_{k=0}^{n}$ is strictly increasing and satisfies $z_{i+1}-z_{i}\geq d_{1}$ for all $i$, we have
$|z_{j}-z_{k}|\geq d_{1}.$
Consequently,
$\left|\frac{2m}{d_{1}}(z_{j}-z_{k})\right|\geq 2m.$
Hence, by the compact support property (M3),
$\sigma\!\left(\frac{2m}{d_{1}}(z_{j}-z_{k})\right)=0,\qquad k\neq j.$
On the other hand,
$\sigma\!\left(\frac{2m}{d_{1}}(z_{j}-z_{j})\right)=\sigma(0)=1.$
Therefore, only the term $k=j$ contributes in both the numerator and the denominator, and we obtain
$\widetilde{T}_{n,r}(g,z_{j})=\frac{\sigma(0)\,P_{r}(g;z_{j},z_{j})}{\sigma(0)}=g(z_{j}).$
This proves the interpolation property. ∎
Theorem 2.7 (Jackson-type approximation estimate).
Let $g\in C^{r+1}[c,d]$. Then there exists a constant $C_{r}>0$, depending only on $r$, such that
$\|\widetilde{T}_{n,r}(g)-g\|_{\infty}\leq C_{r}\,\omega_{r+1}(g,d_{2}),$
where
$d_{2}:=\max_{0\leq k\leq n-1}(z_{k+1}-z_{k})$
denotes the mesh norm of the irregular grids. In particular, if $d_{2}\to 0$, then $\widetilde{T}_{n,r}(g)\to g$ uniformly on $[c,d]$.
Proof.
Let $g\in C^{r+1}[c,d]$ and fix an arbitrary point $y\in[c,d]$. By the definition of the Taylor-accelerated neural network interpolation operator, we have
$\widetilde{T}_{n,r}(g,y)-g(y)=\frac{\sum\limits_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)\big(P_{r}(g;y,z_{k})-g(y)\big)}{\sum\limits_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)}.$
Taking absolute values and using the property (M1), we obtain
$|\widetilde{T}_{n,r}(g,y)-g(y)|\leq\frac{\sum\limits_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)\,|P_{r}(g;y,z_{k})-g(y)|}{\sum\limits_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)}.$ (2.3)
By the compact support property (M3) of $\sigma$, the summands are nonzero only for indices $k$ such that
$\left|\frac{2m}{d_{1}}(y-z_{k})\right|<2m,$
which implies
$|y-z_{k}|<d_{1}\leq d_{2}.$
Fix such an index $k$. Since $g\in C^{r+1}[c,d]$, Taylor’s theorem with integral remainder yields
$|g(y)-P_{r}(g;y,z_{k})|\leq\frac{|y-z_{k}|^{r+1}}{(r+1)!}\,\|g^{(r+1)}\|_{\infty}.$
Fix such an index $k$. Since $g\in C^{r+1}[c,d]$, a classical estimate relating Taylor remainders to the modulus of smoothness (see, e.g., [12, 19]) yields
$|g(y)-P_{r}(g;y,z_{k})|\leq C_{r}\,\omega_{r+1}(g,|y-z_{k}|),$
where $C_{r}>0$ depends only on $r$.
Since $\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)\neq 0$ implies $|y-z_{k}|<d_{1}\leq d_{2}$, and the modulus of smoothness is non-decreasing, we obtain
$|g(y)-P_{r}(g;y,z_{k})|\leq C_{r}\,\omega_{r+1}(g,d_{2}).$
Substituting this estimate into (2.3), we obtain
$|\widetilde{T}_{n,r}(g,y)-g(y)|\leq C_{r}\,\omega_{r+1}(g,d_{2})\,\frac{\sum\limits_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)}{\sum\limits_{k=0}^{n}\sigma\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)}.$
Since the same positive weights appear in the numerator and the denominator, the above fraction equals $1$. Hence,
$|\widetilde{T}_{n,r}(g,y)-g(y)|\leq C_{r}\,\omega_{r+1}(g,d_{2}),\qquad y\in[c,d].$
Taking the supremum over $y\in[c,d]$ completes the proof. ∎
Corollary 2.8.
Let $g\in C^{r+1}[c,d]$ and let
$h_{n}:=\max_{0\leq k\leq n-1}(z_{k+1}-z_{k})$
denote the mesh norm of the irregular grids. Then the Taylor-accelerated neural network interpolation operator $\widetilde{T}_{n,r}$ satisfies
$\|\widetilde{T}_{n,r}(g)-g\|_{\infty}\leq\widehat{C}_{r}\,h_{n}^{\,r+1}\,\|g^{(r+1)}\|_{\infty},$
where $\widehat{C}_{r}>0$ is a constant independent of $n$, depending only on $r$.
Proof.
By Theorem 2.7, there exists a constant $A_{r}>0$, depending only on $r$, such that
$\|\widetilde{T}_{n,r}(g)-g\|_{\infty}\leq A_{r}\,\omega_{r+1}(g,h_{n}).$
Since $g\in C^{r+1}[c,d]$, the classical inequality for higher-order moduli of smoothness (see Section 1) yield
$\omega_{r+1}(g,\delta)\leq B_{r}\,\delta^{\,r+1}\,\|g^{(r+1)}\|_{\infty},\qquad\delta>0,$
where $B_{r}>0$ depends only on $r$.
Applying this estimate with $\delta=h_{n}$, we obtain
$\omega_{r+1}(g,h_{n})\leq B_{r}\,h_{n}^{\,r+1}\,\|g^{(r+1)}\|_{\infty}.$
Combining the above inequalities, we conclude that
$\|\widetilde{T}_{n,r}(g)-g\|_{\infty}\leq(A_{r}B_{r})\,h_{n}^{\,r+1}\,\|g^{(r+1)}\|_{\infty}.$
Setting $\widehat{C}_{r}:=A_{r}B_{r}$ completes the proof. ∎
Corollary 2.9.
Assume that the sequence of irregular grids satisfies the refinement condition: there exists a constant $C>0$, independent of $n$, such that
$h_{n}:=\max_{0\leq k\leq n-1}(z_{k+1}-z_{k})\leq\frac{C}{n}.$
Such a condition is fulfilled, for instance, by any quasi-uniform family of grids of $[c,d]$.
Let $g\in C^{r+1}[c,d]$. Then the Taylor-accelerated neural network interpolation operator $\widetilde{T}_{n,r}$ satisfies
$\|\widetilde{T}_{n,r}(g)-g\|_{\infty}\leq\widetilde{C}_{r}\,n^{-(r+1)}\,\|g^{(r+1)}\|_{\infty},$
where $\widetilde{C}_{r}>0$ is independent of $n$ and depends only on $r$ and the operator parameters.
Proof.
By Corollary 2.8, we have
$\|\widetilde{T}_{n,r}(g)-g\|_{\infty}\leq\widehat{C}_{r}\,h_{n}^{\,r+1}\,\|g^{(r+1)}\|_{\infty}.$
Using the refinement assumption $h_{n}\leq\frac{C}{n}$, it follows that
$h_{n}^{\,r+1}\leq\left(\frac{C}{n}\right)^{r+1}=C^{r+1}\,n^{-(r+1)}.$
Substituting this into the previous inequality yields
$\|\widetilde{T}_{n,r}(g)-g\|_{\infty}\leq\widehat{C}_{r}\,C^{r+1}\,n^{-(r+1)}\,\|g^{(r+1)}\|_{\infty}.$
Defining
$\widetilde{C}_{r}:=\widehat{C}_{r}\,C^{r+1}$
completes the proof. ∎
Remark 2.10 (Optimal order of convergence).
Corollary 2.8 shows that, for functions $g\in C^{r+1}[c,d]$, the Taylor-accelerated neural network interpolation operator $\widetilde{T}_{n,r}$ achieves the optimal approximation order $O(h_{n}^{\,r+1})$, which coincides with the best possible rate for polynomial-based interpolation schemes of degree $r$.
Although the operator introduced in [7] is defined for all continuous functions $g\in C[c,d]$, its higher-order accuracy is obtained through a Lagrange-type reconstruction involving $(r+1)$ neighboring sampling nodes. On irregular grids, this may lead to additional numerical complexity and reduced stability due to the behaviour of Lagrange basis polynomials.
In contrast, the Taylor-accelerated operator proposed in this paper is specifically designed for smooth functions in $C^{r+1}[c,d]$, where local Taylor expansions can be exploited directly. In next Section Numerical experiments confirm that, for sufficiently regular functions, the proposed method produces a significantly faster decay of the approximation error compared with the Lagrange-based interpolation operator of [7]. This highlights the effectiveness of the Taylor acceleration mechanism in practical computations.
2.2 Multivariate Extension
We now extend the above construction to the multivariate setting. Let $d\in\mathbb{N}$ and consider the $d$-dimensional box domain
$\Omega:=[a_{1},b_{1}]\times\cdots\times[a_{d},b_{d}]\subset\mathbb{R}^{d}.$
Let $\eta\in D(m)$ and let $\sigma:\mathbb{R}\to\mathbb{R}$ be the associated activation function introduced in Section 1. We define the multidimensional activation function
$\Psi_{\sigma}:\mathbb{R}^{d}\to\mathbb{R}$
by
$\Psi_{\sigma}(y):=\prod_{i=1}^{d}\sigma(y_{i}),\qquad y=(y_{1},\dots,y_{d})\in\mathbb{R}^{d}.$
The function $\Psi_{\sigma}$ is nonnegative, compactly supported, and satisfies
$\|\Psi_{\sigma}\|_{\infty}:=\sup_{y\in\Omega}|\Psi_{\sigma}(y)|\leq 1.$
Multivariate irregular grids. For each coordinate direction $i\in\{1,\dots,d\}$, we consider an irregular partition of the interval $a_{i},b_{i}]$ given by
$a_{i}=z_{i0}<z_{i1}<\cdots<z_{in}=b_{i}.$
We assume that there exist constants $0<d_{1}\leq d_{2}$, with $d_{2}<2d_{1}$, such that
$d_{1}\leq z_{i,k+1}-z_{ik}\leq d_{2},\qquad k=0,1,\dots,n-1,\ \ i=1,\dots,d.$
For a multi-index $k=(k_{1},\dots,k_{d})\in\{0,\dots,n\}^{d}$, we define the corresponding grid node
$z_{k}:=(z_{1k_{1}},\dots,z_{dk_{d}})\in\Omega.$
As in the one-dimensional case, to ensure consistency near the boundary and preserve the interpolation property, the nodes are extended by setting
$z_{i,n+\ell}:=z_{i,n-r+\ell-1},\qquad\ell=1,\dots,r,\ \ i=1,\dots,d.$
Multivariate Taylor polynomial. Let $g\in C^{r}(\Omega)$. For each node $z_{k}$, we denote by $P_{r}(g;y,z_{k})$ the multivariate Taylor polynomial of total degree $r$ of $g$ centered at $z_{k}$, defined by
$P_{r}(g;y,z_{k})=\sum_{|\alpha|\leq r}\frac{D^{\alpha}g(z_{k})}{\alpha!}(y-z_{k})^{\alpha},\qquad y\in\Omega,$
where $\alpha=(\alpha_{1},\dots,\alpha_{d})\in\mathbb{N}_{0}^{d}$ is a multi-index, $|\alpha|=\alpha_{1}+\cdots+\alpha_{d}$, $\alpha!=\alpha_{1}!\cdots\alpha_{d}!$, and
$D^{\alpha}g:=\frac{\partial^{|\alpha|}g}{\partial y_{1}^{\alpha_{1}}\cdots\partial y_{d}^{\alpha_{d}}}.$
Definition 2.11.
Let $g\in C^{r}(\Omega)$. Then multivariate Taylor-accelerated neural network interpolation operator is defined by
$\widetilde{T}_{n,r}^{(d)}(g,y)=\frac{\displaystyle\sum_{k\in\{0,\dots,n\}^{d}}\Psi_{\sigma}\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)P_{r}(g;y,z_{k})}{\displaystyle\sum_{k\in\{0,\dots,n\}^{d}}\Psi_{\sigma}\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)},\qquad y\in\Omega.$
The extension of the above construction and approximation results from the univariate setting to $\mathbb{R}^{d}$ does not introduce additional analytical difficulties. Since $\mathbb{R}^{d}$ is finite dimensional, all norms are equivalent and the properties of the activation function $\Psi_{\sigma}$ are preserved under the product structure. Consequently, the proofs of convergence, interpolation, and quantitative error estimates follow the same arguments as in the one-dimensional case, with only notational changes and dimension-dependent constants. This situation stands in sharp contrast with approximation problems posed in infinite-dimensional spaces, where genuinely new analytical phenomena typically arise.
3 Numerical Validation
In this section, we present numerical experiments to validate the theoretical approximation results established in the previous sections. In particular, we illustrate the higher-order convergence behaviour of the proposed Taylor-accelerated neural network interpolation operators and provide a direct comparison with the Lagrange-based neural network interpolation operator on irregular grids recently introduced by Costarelli et al. [7].
We consider the smooth target function
$f(y)=\sin(2\pi y)+y^{3},\qquad y\in[0,1],$
which is infinitely differentiable on $[0,1]$ and therefore satisfies the regularity assumptions required in our convergence theorems.
3.1 Activation Function and Grid Construction
Both interpolation operators are based on a compactly supported sigmoidal activation function. In the numerical experiments, we adopt a truncated ramp-type sigmoid $\eta\in D(m)$ defined by
$\eta(y)=\begin{cases}0,&y\leq-m,\\[6.0pt] \dfrac{y+m}{2m},&-m<y<m,\\[10.0pt] 1,&y\geq m,\end{cases}$
where $m>0$ is a fixed support parameter. The associated activation function is given by
$\sigma(y)=\eta(y+m)-\eta(y-m),$
which satisfies $\sigma(0)=1$ and has compact support contained in [2m,2m][-2m,2m].
All experiments were implemented in MATLAB. The sampling nodes were chosen as nested quasi-uniform irregular grids of $[0,1]$, generated by applying small random perturbations to a uniform grid while preserving the mesh regularity condition $\frac{d_{2}}{d_{1}}<2.$ This construction ensures that the theoretical assumptions are satisfied and allows us to study convergence as the number of nodes $n$ increases.
3.2 Error Convergence and Comparison
For each grid size $n$ and Taylor order $r$, we compute the uniform approximation error
$E^{\mathrm{Tay}}_{n,r}:=\bigl\|\widetilde{T}_{n,r}(f)-f\bigr\|_{\infty},$
where $\widetilde{T}_{n,r}$ denotes the proposed Taylor-accelerated operator. In addition, we compute the corresponding error
$E^{\mathrm{Lag}}_{n,r}:=\bigl\|T^{\mathrm{Lag}}_{n,r}(f)-f\bigr\|_{\infty},$
where $T^{\mathrm{Lag}}_{n,r}$ is the Lagrange-based neural network interpolation operator on irregular grids introduced by Costarelli et al.[7].
Figure 1 shows the decay of $E^{\mathrm{Tay}}_{n,r}$ for $r=1,2,3$. As expected, increasing the Taylor order produces a substantial improvement in accuracy, and the error decays faster for larger values of $r$, in agreement with the theoretical rate
$\bigl\|\widetilde{T}_{n,r}(f)-f\bigr\|_{\infty}\leq C\,n^{-(r+1)}.$
Figure 1: Uniform approximation error $E^{\mathrm{Tay}}_{n,r}=\|\widetilde{T}_{n,r}(f)-f\|_{\infty}$ as a function of $n$ for Taylor degrees $r=1,2,3$ on nested quasi-uniform irregular grids. Higher values of $r$ yield significantly improved convergence rates.
To highlight the improvement with respect to the Lagrange-based construction, Figures 2-4 compare the Taylor and Lagrange interpolation operators for each fixed order $r=1,2,3$. In all cases, the Taylor-accelerated interpolation operator provides smaller errors, demonstrating the effectiveness of the local Taylor enrichment.
Figure 2: Error comparison between Taylor and Lagrange neural network interpolation operators for $r=1$.
Figure 3: Error comparison between Taylor and Lagrange neural network interpolation operators for $r=2$.
Figure 4: Error comparison between Taylor and Lagrange neural network interpolation operators for $r=3$.
3.3 Discussion
The numerical results confirm the theoretical findings of the previous sections. The proposed Taylor-accelerated neural network interpolation operators achieve higher-order approximation rates on quasi-uniform irregular grids, and the convergence order increases with the Taylor degree $r$. Moreover, the comparative experiments clearly indicate that the Taylor-accelerated construction provides improved convergence and accuracy with respect to the Lagrange-based irregular grid operator of Costarelli et al.[7], demonstrating the advantages of incorporating local Taylor information into neural network interpolation operators.
4 Conclusion
In this paper, we introduced a new class of Taylor-accelerated neural network interpolation operators defined on quasi-uniform irregular grids. By enriching sampling-based neural network interpolation with local Taylor polynomials, the proposed construction effectively exploits higher-order smoothness information of the target function.
We proved that the operators are well defined, uniformly bounded, and preserve exact interpolation at the grid nodes. Moreover, they reproduce polynomials up to the prescribed order and satisfy direct approximation estimates in terms of higher-order moduli of smoothness, leading to convergence rates of order $r+1$ for sufficiently smooth functions. In particular, under asymptotically quasi-uniform grids, the error decays as
$\|\widetilde{T}_{n,r}(g)-g\|_{\infty}=O\bigl(n^{-(r+1)}\bigr),\qquad n\to\infty.$
Numerical experiments confirmed the theoretical findings and showed that the Taylor-accelerated operator provides improved accuracy compared not only with classical first-order neural network interpolation operators, but also with the recently proposed Lagrange-based neural network interpolation operator on irregular grids. This highlights the advantage of incorporating local Taylor information in achieving higher-order convergence on non-uniform sampling nodes.
Future research directions include adaptive irregular meshes and the development of stochastic or fuzzy variants of the proposed Taylor-accelerated framework.
Declaration of competing interest
The author declares that he has no competing financial interests or personal relationships that could influence the reported work in this paper.